For a data set of the pulse rates for a sample of adult​ females, the lowest pulse rate is 36 beats per​ minute, the mean of the listed pulse rates is x=71.0 beats per​ minute, and their standard deviation is s=12.3 beats per minute.


1
Expert's answer
2021-09-29T00:07:01-0400

Solution:

"\mu=71,\sigma=12.3"

"X\sim N(\mu,\sigma)"

As data is missing, we assume that we need to find "P(X>36)".

"P(X>36)=1-P(X\le 36)
\\=1-P(z\le \dfrac{36-71}{12.3})
\\=1-P(z\le-2.85)
\\=P(z<2.85)
\\=0.99781"